A 32.4 L sealed cylinder contains 120. moles of helium gas at 44.8 atm. What iis its temperature (in Celsius)?
1
Expert's answer
2020-03-26T11:59:35-0400

Solution:

When using the van der Waals equation


"(p+ \frac {a\times v^2}{V^2})(V-b\times v)=v\times R \times T"

we use the following values, taking into account that helium is in the cylinder


"p=44.8\times 101325.011=4.54\times 10^6 \frac {N}{m^2}"

"a=3.38 \times 10^{-3} \frac {N\times m^4}{mole^2}"


"v=120 mole"


"V=3.24\times 10^{-2} m^3"


"b=2.36 \times 10^{-5} \frac {m^3}{mole}"

"R=8.314 \frac {Dg}{mole \times K}"


"T= \frac{(p+\frac {av^2}{V^2}) (V-bv)}{vR}"


"T=136 K"

"t=T-273=-137 C"

Answer: -137 (in Celsius)